OpenStudy (anonymous):

I'll fan you and give you a metal and be your bestest friend if you can answer these 3 questions for me ;D

OpenStudy (anonymous):

What questions?

OpenStudy (anonymous):

Oh crap sin and cos i can't do i am bad at it sorry :(

OpenStudy (anonymous):

1) find the angle by using arctan for 2) x=11*cos22

OpenStudy (anonymous):

3) x=arcsin(7/9), arc = inverse

OpenStudy (anonymous):

4) is Pythagoras them x = sqrt(17^2+15^2)

OpenStudy (anonymous):

2) is cos =adj/hyp Cos22=x/11 cos22x11=x x=10.19 = 10.2

OpenStudy (anonymous):

4) is pythagoras; b^2 = c^2-a^2; then square root your answer

OpenStudy (anonymous):

It's medal, not metal. Also, you might learn something if you read some notes instead of having other people do your homework.

OpenStudy (anonymous):

@Effoff efff the eff offf